I could really use some help. Factor the quadratic expression completely, and find the roots of the expression.Thnak you!
49x2 - 14x - 3

Answer by checkley71(8403)

49X~2-14X-3=0 OR (7X-3)(7X+1)=0 OR 7X-3=0 OR 7X=3 OR X=3/7 & 7X+1=0 OR 7X=-1 OR X=-1/7